OpenStudy (stranger.tz):

QN1) 5 cow and 20 goats can graze a given piece of pasture land for 16 days. if 5 more cows are purchased by the owner, for how long grazing pasture will suffice? QN2) How long will take for 6 cows and 20 goats to graze in the land that 6 cows or 20 goats graze a piece of pasture for 16 days?

OpenStudy (anonymous):

(5 x +20 y) = 16 days grazing rate for goats and cows right?

OpenStudy (stranger.tz):

what is the Ans??

OpenStudy (mrnood):

you cannot answer this with the information available You have 2 unknowns (rate of grazing for cows and goats) and only 1 equation

OpenStudy (mrnood):

also - you should not expect or demand 'the answer'. The purpose of this forum is ot guide you to get the answer for yourself.

OpenStudy (mrnood):

@3abf6277 I think that that is not quite the correct equation. There is a fixed amount of pasture available - let's call it P if cows eat x /day and goats eat y per day then we know 16(5x + 20y) = P Adding 5 cows gives Z(10x+20y)=P (where Z is number of days that pasture will provide for 5 more cows) so Z(10x+20y)=16(5x + 20y)

OpenStudy (stranger.tz):

THANK YOU Friends